Y^2+x^2p^2-2xyp=4/p^2find GS and SS
Where GS=y^2-2cxy+c^2(x^2-1)=m^2
SS= y^2+m^2=m^2 answers to be got
1
Expert's answer
2020-10-26T20:37:37-0400

Given equation is incorrect

Correct equation is,

"y^2+x^2p^2-2xyp=4"


"(y-xp)^2=2^2"


"\to y-xp=2"


"\to y-x\dfrac{dy}{dx}=2"


"\to y-2=x\dfrac{dy}{dx}"


"\to \dfrac{dy}{y-2}=\dfrac{dx}{x}"

Integrating Both the sides

"\to \int \dfrac{dy}{y-2}=\int\dfrac{dx}{x}" '


"\therefore log(y-2)=logx+logc"


"\to y-2=xc"


"\to y=xc+2"


For general solution ,

as y=xc+2


or y-xc=2


Squaring on both sides

"\to (y-xc)^2=2^2"


"\to y^2+x^2c^2-2xyc=4\\"

Subtract "c^2" on both sides

"\to y^2+x^2c^2-2xyc-c^2=4-c^2\\
\to y^2-2xyc+c^2(x^2-1)=m^2" ( where "m^2=4-c^2" )


For SS


Let calculate value of constant at (0,0)

"\to 0+c^2(-1)=m^2\\
c^2=-m^2"


So SS is

"\to y-0+c^2(-1)=m^2\\
\to y-c^2=m^2\\
\to y+m^2=m^2"